OpenStudy (anonymous):

Find the quotient z/w if z=12[cos(5pi/18) +isin(5pi/18)] and w=4[cos(7pi/36) +isin(7pi/36)]

OpenStudy (anonymous):

why dont you first transform them into the form \[r*e^{(i*\Theta)}\]

OpenStudy (anonymous):

then you would have fraction of exponents so you would just substract the exponents

OpenStudy (loser66):

@Lills knock,knock understand what he means?

OpenStudy (anonymous):

My Internet went out, but I'm not exactly sure what he means

OpenStudy (anonymous):

subtract the angles

OpenStudy (anonymous):

\(12\div 4=3\) and \(\frac{5\pi}{18}-\frac{7\pi}{36}\) will give you the angle you need

OpenStudy (anonymous):

that is why this easy and not hard it is just a subtraction problem

OpenStudy (loser66):

but the condition is student know what the Euler number is, hehehe. me not for example

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I'm not quite sure i understand

OpenStudy (loser66):

hahahaha..... that can be foreseen.

OpenStudy (anonymous):

in plain english subtract the angles when you divide

OpenStudy (anonymous):

Like how to write the final answer

OpenStudy (anonymous):

\[r_1\left(\cos(\theta)+i\sin(\theta)\right)\div r_2\left (\cos(\alpha)+i\sin(\alpha)\right)\] \[=\frac{r_1}{r_2}\left (\cos(\theta-\alpha)+i\sin(\theta-\alpha)\right)\]

OpenStudy (anonymous):

in your case \(12\div 4=3\) so that goes out front

OpenStudy (anonymous):

your next job is only to compute \[\frac{5\pi}{18}-\frac{7\pi}{36}\]

OpenStudy (anonymous):

i totaly forgot about that formula , i guess hes supposed to use that formula not convert

OpenStudy (anonymous):

right it is just a subtraction problem that is all

OpenStudy (loser66):

Please, first off, the asker has to know about Euler formula. what if he/she didn't know about it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

divide the modulus subtract the anglers that is all

OpenStudy (anonymous):

So the final answer would be 3(cos(pi/12)+isin(pi/12) ?

OpenStudy (anonymous):

yes, that is it

OpenStudy (anonymous):

easy right? boils down only to dividing 12 by 4 and subtraction that is why you write complex numbers in this form makes the calculation really really easy

OpenStudy (anonymous):

Yes thank you so much for your help tonight!